Circular No. 8576

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1188 SCORPII = NOVA SCORPII 2005 N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

informs us that the designation V1188 Sco has been given to this nova (cf. IAUC 8574, 8575).

H. Naito and N. Tokimasa, Nishi-Harima Astronomical Observatory (NHAO); and H. Yamaoka, Kyushu University, report that a low-resolution spectrum (range 300–850 nm; resolution 150) of V1188 Sco, obtained on July 27.6 UT with NHAO 0.60-m telescope (+ NILS), shows a broad Hα feature (FWHM = 1730 km/s) with P-Cyg profile, suggesting that V1188 Sco is indeed a classical nova. Yamaoka adds that M. Fujii, Ibara, Japan, obtained a spectrum on July 27.53, which shows a clear Hα line in emission.

F. M. Walter, State University of New York at Stony Brook; H. E. Bond, Space Telescope Science Institute; and A. Pasten, SMARTS and Cerro Tololo Interamerican Observatory (CTIO), confirm that V1188 Sco is a classical nova near or before peak. A 0.31-nm-resolution spectrum (range 570–690 nm), taken on July 28.16 UT using the SMARTS/CTIO 1.5-m RC spectrograph, shows broad Hα (FWZI = 4000 km/s, EW= −3.5 nm) and Fe II (74) lines in emission atop a photospheric continuum with absorption lines. He I 587.6-nm appears in absorption; the Na D line (EW = 0.5 nm) shows three absorption components. The peak of the Hα emission is blueshifted by ∼ 250 km/s; the P-Cyg absorption is blueshifted by ∼ 900 km/s. There may also be an absorption feature at +350 km/s.

COMETS C/2005 M6–M10 (SOHO) Several more tailless Kreutz sungrazing comets have been found on

SOHO images (cf. IAUC 8572). C/2005 M6 was tiny, faint, and diffuse. C/2005 M7 was stellar, tiny, and faint in the C3 images, but slightly diffuse in C2 images, reaching mag ∼ 6.4 at 7.8R_{⊙} on June 26.021 UT. C/2005 M8 was faint and diffuse, reaching mag ∼ 8.0 at 7.6R_{⊙} on June 27.829. C/2005 M9 was very diffuse in C2 images, reaching mag ∼ 7.8 at 7.5R_{⊙} on June 28.162. C/2005 M10 was diffuse and very faint.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 M6 June 24.777 6^{h}08.^{m}4 +21^{o}40′ C2  SH 2005-O35
C/2005 M7  25.529 6 10.1 +20 29 C3/2 XL 2005-O35
C/2005 M8 27.771 6 19.5 +21 41 C2 RM 2005-O35
C/2005 M9  27.846 6 19.2 +20 55 C3/2 RM 2005-O35
C/2005 M10  29.421 6 26.4 +21 38 C2  KC 2005-O35
2005 July 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green